# Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation

### v1

@@ 4.5.1 Component-wise Analysis
We evaluate the necessity of each core component by systematically removing individual modules from the complete Self-Foveate framework.
Table 6 presents the ablation results on the SQuAD dataset using Llama-3.1-8B as the base model. The results demonstrate that each component contributes significantly to the overall performance.
Setting Recall LLM Acc.
w/o Micro-Foveate 0.283 0.277
w/o Scatter-Foveate 0.274 0.260
w/o Macro-Foveate 0.344 0.339
Self-Foveate (Full) 0.484 0.490
@@
The removal of any single component leads to performance degradation, validating the complementary necessity of all three core components.
@@ 4.5.2 Answer Regeneration Analysis
In our reverse synthesis paradigm, although foveate elements (e.g., "QuantumX processor") can directly serve as answers, we regenerate answers to enhance fluency, completeness, and semantic coherence. To validate this design choice, we compare the quality of answers with and without the regeneration mechanism.
Table 7 presents the comparative analysis on the SQuAD dataset, evaluating both fluency and completeness of generated answers through GPT-4o evaluation. The results demonstrate significant improvements in both aspects when employing answer regeneration.
Setting High (%) Medium (%) Low (%)
Fluency
w/o Regeneration 64.9 19.5 15.6
Reverse Synthesis 93.0 6.8 0.2
Completeness
w/o Regeneration 36.7 27.5 35.7
Reverse Synthesis 76.6 19.7 3.7
@@
The results reveal that answer regeneration dramatically improves both fluency and completeness. This validates our methodological choice to maintain consistency across all synthesis paradigms by generating contextually appropriate and semantically coherent answers from the source text, rather than directly using extracted foveate elements.
@@ 5 Conclusion
In this paper, we proposed Self-Foveate, an unsupervised text-based instruction generation method powered by LLMs, which further explores key challenges in diversity and difficulty. Its innovative multi-level foveation and Re-synthesis Module ensure the acquisition of high-quality, text-faithful instructions from unsupervised text. Extensive experiments, including performance evaluations on downstream tasks, demonstrate the effectiveness of Self-Foveate. Our study has obtained many insightful findings, laying the foundation for future research in areas such as instruction difficulty assessment.
@@ Acknowledgments
Although Self-Foveate provides significant improvements in generating diverse and difficulty-oriented fine-tuning instruction data based on unsupervised text data, several limitations must be acknowledged. First, the computational cost of processing large-scale unsupervised text data using closed-source SOTA large language models (LLMs) remains substantial. Ensuring data quality requires multi-step reasoning and iterative processing on unsupervised text data, which further escalates computational demands. Second, although Self-Foveate aims to fully synthesize data from unsupervised text and incorporates verification mechanisms, the inherent tendency of LLMs to hallucinate or generate erroneous information persists as a challenge. While precise prompting strategies and higher-quality unsupervised text data can mitigate these inaccuracies, they cannot be entirely eliminated.
@@
We state that any research or application arising from this study is strictly authorized solely for research purposes. In our work, any unsupervised text datasets used are from public sources and do not contain any private information. In this paper, we have fully presented the prompts used by Self-Foveate in the Appendix. All synthesized instructions rely on the provided unsupervised text and are inspected by relevant modules. Therefore, our method strives to minimize potential safety and ethical risks as much as possible. However, during the process of synthesizing fine-tuning instruction data, maliciously provided unsupervised text data can lead the model to produce harmful or inappropriate outputs, which is a shared problem. Additionally, potential unfairness and discrimination present in the unsupervised text data might be amplified by LLMs during the instruction generation process. Ensuring the quality of generated fine-tuning instruction data in a safe and highly controllable manner is crucial. The application of these techniques should be guided by ethical considerations, with safeguards in place to prevent misuse and reduce the likelihood of producing harmful outcomes.
@@ A.2 Usage of Datasets
During the synthesis process, we use the articles or contexts from the dataset (excluding the questions) and generate instructions through the Self-Foveate. In the evaluation of problem-solving capabilities, we assess the fine-tuned models using the corresponding questions in the dataset.
Appendix B Instruction Synthesis
@@
Taking into account both performance and API costs, and to eliminate potential biases in experimental results caused by specific teacher models (such as prompt hacking), we employ two models, GPT-4o mini(gpt-4o-mini-2024-07-18)(OpenAI, 2024) and DeepSeek-V3(DeepSeek-AI, 2024), to generate instructions from unsupervised text using the same methodology and sampling hyperparameters. Detailed hyperparameters are provided in Table 8, where, as described in Section 3, the normal mode is utilized for the initial foveate synthesis, and the high-creativity mode is adopted for re-synthesis.
Hyperparameters Values
Normal Mode
frequency_penalty 0.5
max_completion_tokens None
presence_penalty 0
temperature 0.5
top_p 1.0
High-creativity Mode
frequency_penalty 0.5
max_completion_tokens None
presence_penalty 0
temperature 1.2
top_p 1.0
Table 8: Hyperparameters of Synthetic Task Generation.
Appendix C Diversity Experiment

### v2

@@ 4.5.1 Component-wise Analysis
We evaluate the necessity of each core component by systematically removing individual components from the complete Self-Foveate framework. Table 6 presents the ablation results on the SQuAD dataset using Llama-3.1-8B as the base model.
Setting Recall LLM Acc. w/o Micro-Foveate 0.283 0.277 w/o Scatter-Foveate 0.274 0.260 w/o Macro-Foveate 0.344 0.339 Self-Foveate (Full) 0.484 0.490
@@
The removal of any single component leads to clear performance degradation, with scatter-foveate removal causing the largest drop (Recall from 0.484 to 0.274), confirming that all three components are necessary and complementary.
@@ 4.5.2 Answer Regeneration Analysis
In our reverse synthesis paradigm, although foveate elements (e.g., “QuantumX processor”) can directly serve as answers, we regenerate answers to enhance fluency, completeness, and semantic coherence. To validate this design choice, we compare the quality of answers with and without the regeneration mechanism. Table 7 presents the comparative analysis on the SQuAD dataset, evaluating both fluency and completeness of synthesized answers using GPT-4o.
Setting High (%) Medium (%) Low (%) Fluency w/o Regeneration 64.9 19.5 15.6 Reverse Synthesis 93.0 6.8 0.2 Completeness w/o Regeneration 36.7 27.5 35.7 Reverse Synthesis 76.6 19.7 3.7
@@
The results show that answer regeneration substantially improves both fluency and completeness. This validates our choice to maintain consistency across all synthesis paradigms by synthesizing contextually appropriate and semantically coherent answers from the source text, rather than directly using extracted foveate elements.
@@ 5 Conclusion
In this paper, we presented Self-Foveate, an LLM-driven methodology for synthesizing diverse and challenging instructions from unsupervised text. Drawing inspiration from hierarchical human visual perception, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology: micro-foveate captures fine-grained entity details, scatter-foveate integrates dispersed cross-entity relationships, and macro-foveate extracts holistic rhetorical patterns. A re-synthesis module further enhances fidelity to source text and overall instruction quality. Comprehensive experiments across multiple unsupervised corpora and diverse model architectures demonstrate that Self-Foveate consistently outperforms existing methods in instruction diversity, difficulty, and downstream task performance. Our work establishes multi-level foveation as an effective paradigm for high-quality instruction synthesis from unsupervised text.
@@ Acknowledgments
Although Self-Foveate provides significant improvements in synthesizing diverse and difficulty-oriented fine-tuning instruction data based on unsupervised text, several limitations must be acknowledged. First, the computational cost of processing large-scale unsupervised text using closed-source state-of-the-art large language models (LLMs) remains substantial. Ensuring data quality requires multi-step reasoning and iterative processing on unsupervised text, which further escalates computational demands. Second, although Self-Foveate aims to fully synthesize data from unsupervised text and incorporates verification mechanisms, the inherent tendency of LLMs to hallucinate or generate erroneous information persists as a challenge. While precise prompting strategies and higher-quality unsupervised text can mitigate these inaccuracies, they cannot be entirely eliminated.
@@
We state that any research or application arising from this study is strictly authorized solely for research purposes. In our work, any unsupervised text datasets used are from public sources and do not contain any private information. In this paper, we have fully presented the prompts used by Self-Foveate in the Appendix. All synthesized instructions rely on the provided unsupervised text and are inspected by relevant modules. Therefore, our method strives to minimize potential safety and ethical risks as much as possible. However, during the process of synthesizing fine-tuning instruction data, maliciously provided unsupervised text can lead the model to produce harmful or inappropriate outputs, which is a shared problem. Additionally, potential unfairness and discrimination present in the unsupervised text might be amplified by LLMs during the instruction synthesis process. Ensuring the quality of synthesized fine-tuning instruction data in a safe and highly controllable manner is crucial. The application of these techniques should be guided by ethical considerations, with safeguards in place to prevent misuse and reduce the likelihood of producing harmful outcomes.
@@ A.2 Usage of Datasets
During the synthesis process, we use the articles or contexts from the dataset (excluding the questions) and synthesize instructions through the Self-Foveate. In the evaluation of problem-solving capabilities, we assess the fine-tuned models using the corresponding questions in the dataset.
Appendix B Instruction Synthesis Details
B.1 Teacher Models and Sampling Hyperparameters
@@
Taking into account both performance and API costs, and to eliminate potential biases in experimental results caused by specific teacher models (such as prompt hacking), we employ two models, GPT-4o mini(gpt-4o-mini-2024-07-18)(OpenAI et al., 2024) and DeepSeek-V3(DeepSeek-AI et al., 2025), to synthesize instructions from unsupervised text using the same methodology and sampling hyperparameters. Detailed hyperparameters are provided in Table 8, where, as described in Section 3, the normal mode is utilized for the initial foveate synthesis, and the high-creativity mode is adopted for re-synthesis.
Hyperparameters Values Normal Mode frequency_penalty 0.5 max_completion_tokens None presence_penalty 0 temperature 0.5 top_p 1.0 High-creativity Mode frequency_penalty 0.5 max_completion_tokens None presence_penalty 0 temperature 1.2 top_p 1.0
Table 8: Hyperparameters of Synthetic Instruction Synthesis.
B.2 Adaptive Task Count Determination
To adaptively determine the number of instructions to synthesize for each article, Self-Foveate employs a two-stage process: (1) counting contextually relevant words via semantic similarity thresholding, and (2) normalizing the resulting counts using a Box-Cox transformation. This ensures that articles with richer content receive proportionally more instructions while maintaining a balanced distribution across the dataset.
Context Thresholding.
For each article $d_{i}$, we compute the text-level embedding $\mathbf{e}_{\text{text}}=\text{Embed}(d_{i})$ and the word-level embeddings $\mathbf{e}_{w_{j}}=\text{Embed}(w_{j})$ for each unique word $w_{j}$ in $d_{i}$. A word is considered contextually relevant if its cosine similarity with the full text exceeds the threshold $\tau$:
$c_{i}=\left|\left\{w_{j}:\text{CosSim}(\mathbf{e}_{\text{text}},\mathbf{e}_{w_{j}})\geq\tau\right\}\right|$ (2)
where $\tau=0.16$ (CONTEXT_THRESHOLD) by default.
Box-Cox Normalization.
The raw counts $\{c_{i}\}$ are then normalized using the Box-Cox transformation to produce the final task counts. The procedure is detailed in Algorithm 1.
Algorithm 1 Adaptive Task Count via Box-Cox Normalization
1:Article set $\mathcal{D}=\{d_{1},\ldots,d_{N}\}$
2:   Threshold $\tau=0.16$, target mean $\mu=8.0$
3:   Scaling factor $\alpha=1.0$
4:Task counts $\{n_{i}\}$ for each article $d_{i}$
5:for each article $d_{i}\in\mathcal{D}$ do
6:  $\mathbf{e}_{\text{text}}\leftarrow\text{Embed}(d_{i})$
7:  $W\leftarrow\text{UniqueWords}(d_{i})$
8:  for each word $w_{j}\in W$ do
9:   $\mathbf{e}_{w_{j}}\leftarrow\text{Embed}(w_{j})$
10:   $s_{j}\leftarrow\text{CosSim}(\mathbf{e}_{\text{text}},\mathbf{e}_{w_{j}})$
11:  end for
12:  $c_{i}\leftarrow|\{w_{j}:s_{j}\geq\tau\}|$
13:end for
14:// Ensure positivity for Box-Cox
15:if $\min(\{c_{i}\})\leq 0$ then
16:  $c_{i}\leftarrow c_{i}+|\min(\{c_{i}\})|+1,\quad\forall\,i$
17:end if
[…]
21:return $\{n_{i}\}$
B.3 Scatter Keyword Grouping
The scatter-foveate level synthesizes instructions from keyword groups of varying sizes (1, 2, or 3 keywords per group). Given a target number of scatter-level instructions $n_{\text{scatter}}$ for an article, the number of groups at each size is determined by two configuration ratios, as described in Algorithm 2.
Algorithm 2 Scatter Keyword Grouping
1:Target instruction count $n_{\text{scatter}}$, ratios $r_{a}=2.50$, $r_{b}=2.25$
2:Keyword groups $\mathcal{G}$
3:// Compute group distribution
[…]
7:// Compute total keywords needed
[…]
9:// Extract and rank keywords via LLM
[…]
13:// Construct groups
[…]
17:return $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$
The ratios $r_{a}$ (SCATTER_RATIO_A =2.50=2.50) and $r_{b}$ (SCATTER_RATIO_B =2.25=2.25) are designed to ensure a balanced distribution of keyword group sizes. Larger groups (3 keywords) produce more complex, multi-aspect instructions, while single-keyword groups yield focused, targeted questions. By including important (core and major) keywords into the shuffled pool, we ensure that critical concepts appear across multiple group sizes, further enhancing instruction diversity.
B.4 Re-synthesis Module
The re-synthesis module handles cases where initial instruction generation fails, either due to invalid format or unanswerable content. The process follows a structured retry mechanism:
Validation Criteria.
A generated instruction-answer pair is considered valid if it satisfies all of the following:
  1. 1.
    The output matches the expected “Question: ... Answer: ...” format (verified via regex parsing).
  2. 2.
    The answer does not contain “I don’t know,” indicating the model could not derive an answer from the background knowledge.
  3. 3.
    The instruction and answer are both non-empty after stripping whitespace.
Re-synthesis Flow.
For each failed instruction, the re-synthesis module applies the retry procedure described in Algorithm 3.
Algorithm 3 Re-synthesis Retry Procedure
1:Failed index set $\mathcal{F}$, successful pairs $\mathcal{S}$, article $d$, max retries $R=5$
2:Re-synthesized instruction set $\mathcal{I}_{\text{new}}$
3:for each failed index $i\in\mathcal{F}$ do
4:  $f_{i}\leftarrow\text{GetFeature}(i)$ $\triangleright$ keyword / group / sentence
5:  for $r=1$ to $R$ do
6:   $\mathcal{E}\leftarrow\text{Shuffle}(\mathcal{S})$ $\triangleright$ Few-shot examples
7:   $(q,a)\leftarrow\text{LLM}(d,\;f_{i},\;\mathcal{E})$
8:   if $\text{IsValid}(q,a)$ then $\triangleright$ Format + answerable + non-empty
9:     $\mathcal{I}_{\text{new}}\leftarrow\mathcal{I}_{\text{new}}\cup\{(q,a)\}$
10:     break
11:   end if
12:  end for
13:end for
14:return $\mathcal{I}_{\text{new}}$
B.5 Complete Hyperparameter Summary
Table 9 provides a complete summary of all framework-level hyperparameters used in the Self-Foveate pipeline.
Parameter Value Description Thresholding & Normalization CONTEXT_THRESHOLD 0.16 Cosine similarity threshold for context relevance DEFAULT_TARGET_MEAN ($\mu$) 8.0 Target mean for Box-Cox normalization DEFAULT_ALPHA ($\alpha$) 1.0 Scaling factor for Box-Cox transformation Scatter Keyword Grouping SCATTER_RATIO_A 2.50 Ratio for single-keyword groups SCATTER_RATIO_B 2.25 Ratio for two-keyword groups Processing Configuration DEFAULT_MAX_RETRIES 5 Maximum re-synthesis attempts DEFAULT_MAX_WORKERS 96 Parallel processing threads (articles) EMBEDDING_MAX_WORKERS 10 Parallel threads for embedding generation
Table 9: Complete hyperparameter configuration of the Self-Foveate framework.
Appendix C Prompt Templates – Micro-foveate Level
This section presents all prompt templates used in the Micro-foveate level of the Self-Foveate pipeline. The micro-level strategy extracts individual keywords from the article and synthesizes instructions based on those keywords.
C.1 Keywords Extraction
Micro Keywords Extraction – System Prompt You are a professional text analysis expert. Your task is to extract keywords from the given article. Based on the following article, please extract {num_micro} semantically different keywords in random order, with: Please ensure: 1. All keywords are based on the article content. 2. Each keyword is unique to maintain diversity. 3. Keywords can be single words or meaningful phrases. 4. Keywords should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. keyword1 2. keyword2
Micro Keywords Extraction – User Prompt Article content: {content}
C.2 Instruction Generation
Micro Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_micro} questions/instructions based on the given background knowledge and answers. Based on the following background knowledge and answers, please generate one question/instruction for each answer. Requirements: 1. Each generated question must be answerable by its corresponding answer 2. Questions/instructions must be based on the background knowledge content 3. Each question/instruction should focus on different aspects or information from the background knowledge 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Micro Instruction Generation – User Prompt The background knowledge is: {content} The answers are: {answers} Please generate {num_micro} questions/instructions based on the background knowledge and answers above.
C.3 Instruction Regeneration
Micro Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target answer. Based on the following background knowledge and answers, please generate one question/instruction for the target answer. Requirements: 1. The generated question/instruction must be answerable by its corresponding answer. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Micro Instruction Regeneration – User Prompt The background knowledge is: {content} The target answer is: {answer}
Appendix D Prompt Templates – Scatter-foveate Level
This section presents all prompt templates used in the Scatter-foveate level of the Self-Foveate pipeline. The scatter-level strategy extracts keyword groups of varying sizes (1, 2, or 3 keywords per group) and synthesizes instructions that incorporate all keywords from each group.
D.1 Keywords Extraction
Scatter Keywords Extraction – System Prompt You are a professional text analysis expert. Your task is to extract keywords from the given article. Based on the following article, please extract {num_keywords} semantically different keywords in random order, with: Please ensure: 1. All keywords are based on the article content. 2. Each keyword is unique to maintain diversity. 3. Keywords can be single words or meaningful phrases. 4. Keywords should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. keyword1 2. keyword2
Scatter Keywords Extraction – User Prompt Article content: {content}
D.2 Important Keywords Selection
Scatter Important Keywords Selection – System Prompt You are a professional text analysis expert. Your task is to identify the most important keywords from the given background knowledge. Please analyze the given keywords and select: 1. {num_core} core keywords that are absolutely critical (primary keywords) 2. {num_major} major keywords that are highly important but secondary (secondary keywords) For selecting the core keyword (select exactly {num_core}): - Must be the most essential concept that captures the central theme - Should be impossible to understand the text without this keyword - Removing this keyword would make the text lose its main focus For selecting major keywords (select exactly {num_major}): - Should be key supporting concepts that elaborate on the core keyword - Must be directly relevant to explaining or contextualizing the main topic - Should cover different important aspects of the content Please output strictly in this format without any other text: Core keyword: 1. [first core keyword] 2. [second core keyword] Major keywords: 1. [first major keyword] 2. [second major keyword]
Scatter Important Keywords Selection – User Prompt Background knowledge: {content} Keywords: {keywords}
D.3 Instruction Generation
Scatter Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_scatter} questions/instructions based on the given background knowledge and keyword groups. Based on the following background knowledge and keyword groups, please generate one question/instruction for each keyword group that incorporates ALL keywords from that group. Requirements: 1. Each generated question/instruction must use ALL keywords from its corresponding group 2. Questions/instructions must be based on the background knowledge content 3. Each question/instruction should focus on different aspects or information from the background knowledge 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Scatter Instruction Generation – User Prompt The background knowledge is: {content} The keyword groups are: {keyword_groups} Please generate {num_scatter} questions/instructions based on the background knowledge and keyword groups above.
D.4 Instruction Regeneration
Scatter Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target keyword group that incorporates ALL keywords from that group. Based on the following background knowledge and keyword groups, please generate one question/instruction for the target keyword group that incorporates ALL keywords from that group. Requirements: 1. The generated question/instruction must use ALL keywords from its corresponding group. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Scatter Instruction Regeneration – User Prompt The background knowledge is: {content} The target keyword group to use: {keyword_group}
Appendix E Prompt Templates – Macro-foveate Level
This section presents all prompt templates used in the Macro-foveate level of the Self-Foveate pipeline. The macro-level strategy extracts complete sentences as contextual features and synthesizes instructions based on those sentences.
E.1 Sentence Extraction
Macro Sentence Extraction – System Prompt You are a professional text analysis expert. Your task is to extract sentences from the given article. Based on the following article, please extract {num_macro} semantically different key sentences in random order, with: Please ensure: 1. All key sentences are based on the article content. 2. Each key sentence is unique to maintain diversity. 3. Key sentences should focus on identifying key details, focusing on transitions, quotations, comparisons, and rhetorical devices. 4. Key sentences should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. key sentence1 2. key sentence2
Macro Sentence Extraction – User Prompt Article content: {content}
E.2 Instruction Generation
Macro Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_macro} questions/instructions based on the given background knowledge and key sentences. Based on the following background knowledge and key sentences, please generate one question/instruction for each key sentence. Requirements: 1. Each generated question/instruction must use the corresponding key sentence. 2. Questions/instructions must be based on the background knowledge content. 3. Each question/instruction should focus on different aspects or information from the background knowledge. 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Macro Instruction Generation – User Prompt The background knowledge is: {content} The key sentences are: {sentences} Please generate {num_macro} questions/instructions based on the background knowledge and key sentences above.
E.3 Instruction Regeneration
Macro Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target key sentence. Based on the following background knowledge and key sentences, please generate one question/instruction for the target key sentence. Requirements: 1. The generated question/instruction must use the corresponding key sentence. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Macro Instruction Regeneration – User Prompt The background knowledge is: {content} The target key sentence to use: {sentence}
Appendix F Prompt Templates – Task Generation
This section presents all prompt templates used in the Task Generation module of the Self-Foveate pipeline. This module handles reading comprehension answer generation and answer improvement through rereading.
F.1 Reading Comprehension (Answer Generation)
Reading Comprehension (Answer Generation) – System Prompt You are an excellent reading comprehension assistant designed to help users answer the question based on provided background knowledge. Please carefully read the background knowledge and accurately answer the question posed based on its content. - Do not alter key information from the original text. - Avoid using phrases like “based on the above article.” - If the background knowledge lacks sufficient information to answer a question, respond with “I don’t know.” - Ensure your answers are derived from the background knowledge, detailed, and accurate. - Answer the question thoroughly. Please generate the answer following this format strictly: Question: [question] Answer: [detailed and accurate answer]
Reading Comprehension (Answer Generation) – User Prompt The background knowledge is: {unsupervised_knowledge_data} Please answer the following question based on the content of the article above: {instruction}
F.2 Answer Rereading (Answer Improvement)
Answer Rereading (Answer Improvement) – System Prompt You are a smart AI assistant. For a given question-answer pair, improve the answer by correcting errors, bolstering informativeness, aligning with the question, and providing comprehensive detail. - Do not alter key information from the original text. - Avoid using phrases like “based on the above article.” - Ensure your answers are derived from the background knowledge, detailed, and accurate. - Answer the question thoroughly. Please generate the rewritten answer following this format strictly: Question: [question] Answer: [detailed and accurate rewritten answer]
Answer Rereading (Answer Improvement) – User Prompt The background knowledge is: {unsupervised_knowledge_data} The question is: {instruction} The original answer is: {output} Please reread the background knowledge and generate a more detailed and accurate rewritten answer based on the content of the article above.
Appendix G Diversity Experiment